# One Demo Is All It Takes: Planning Domain Derivation with LLMs from A Single Demonstration

### v1

@@ Abstract
Pre-trained Large Language Models (LLMs) have shown promise in solving planning problems but often struggle to ensure plan correctness, especially for long-horizon tasks. Meanwhile, traditional robotic task and motion planning (TAMP) frameworks address these challenges more reliably by combining high-level symbolic search with low-level motion planning. At the core of TAMP is the planning domain, an abstract world representation defined through symbolic predicates and actions. However, creating these domains typically involves substantial manual effort and domain expertise, limiting generalizability. We introduce Planning Domain Derivation with LLMs (PDDLLM), a novel approach that combines simulated physical interaction with LLM reasoning to improve planning performance. The method reduces reliance on humans by inferring planning domains from a single annotated task-execution demonstration. Unlike prior domain-inference methods that rely on partially predefined or language descriptions of planning domains, PDDLLM constructs domains entirely from scratch and automatically integrates them with low-level motion planning skills, enabling fully automated long-horizon planning. PDDLLM is evaluated on over 1,200 diverse tasks spanning nine environments and benchmarked against six LLM-based planning baselines, demonstrating superior long-horizon planning performance, lower token costs, and successful deployment on multiple physical robot platforms.
@@ 1 Introduction
Robotic planning remains challenging in complex scenarios requiring abstract, long-horizon reasoning. Large language models (LLMs) show strong generalization in this domain but often struggle with temporal dependencies in long-horizon tasks. Task and Motion Planning (TAMP) frameworks offer a solution by integrating high-level symbolic reasoning with low-level motion planning, yet they rely on manually designed planning domains, abstract world models that are labor-intensive to create and limit adaptability. To overcome this, recent advances in world model learning aim to support long-horizon decision making via learned world models (Guan et al., 2023; Zhao et al., 2024; Liang et al., 2024; Tang et al., 2024), using either natural language or formal definitions like planning domain definition language (PDDL) (McDermott et al., 1998). Building on this paradigm, our approach leverages the concept of planning domains from TAMP to enhance the long-horizon reasoning capability of LLMs. By combining LLMs with physical simulation, we automatically generate a task-specific PDDL planning domain from a single expert demonstration and a brief task description. The resulting PDDL planning domain can be seamlessly integrated with a low-level motion planner, enabling the system to solve tasks with greater complexity than the original demonstration.
Our method is the first to construct planning domains entirely from scratch, without relying on any predefined predicates, actions, or detailed human descriptions. Furthermore, it fully automates the deployment of planning domains by automatically integrating them with low-level motion planners, eliminating the need for manual intervention. The main contributions of this paper are as follows:
  • A novel framework that integrates an automatically-generated planning domain with a low-level motion planner to address long-horizon robotic planning tasks.
  • An algorithm that combines LLMs and physical simulation to automatically generate a human-interpretable planning domain from a single human demonstration.
  • We introduce a logic-constrained action sampler (LoCAS), which automatically integrates the derived planning domain with low-level motion planning for effective task execution.
  • We evaluate PDDLLM on over 1,200 tasks across nine environments, demonstrating superior performance and more efficient token usage compared to state-of-the-art baselines. Furthermore, we successfully deploy PDDLLM on two physical robot platforms.
@@ 2 Related Work
A recent line of research aims to infer the planning domain directly from human demonstrations, environment interactions, or natural language. However, these approaches often depend on partially or fully predefined symbolic predicates and actions (Silver et al., 2023; Kumar et al., 2023; Huang et al., 2024). Some recent studies have explored leveraging LMs for domain generation, primarily by extending manually defined domains with additional predicates and actions (Liang et al., 2024; Athalye et al., 2024). The approach proposed by Guan et al. (2023) generates planning domains requiring prompts containing expert-crafted PDDL domain examples and detailed descriptions of predicates and actions. Moreover, many of these studies assume that the robot is already equipped with motion planning skills aligned with the logical actions (Kumar et al., 2023; Huang et al., 2024; Liang et al., 2024; Athalye et al., 2024). This assumption necessitates manual integration between the symbolic planning domain and low-level motion planners, thereby constraining scalability and limiting autonomy.
@@ 3 Preliminaries
PDDL is a standard formal language used to specify planning problems. The object set $\mathcal{O}$ represents the environment’s objects, whose continuous state $\mathcal{S}$, such as pose, color, and size, can be queried via a perception function $\mathcal{I}:\mathcal{O}\times\mathcal{I}\rightarrow\mathcal{S}$. The PDDL domain $\mathcal{D}=(\mathcal{P},\mathcal{A})$ describes the general rules of the environment, consisting of a set of logical predicates $\mathcal{P}$ and a set of logical actions $\mathcal{A}$. A logical predicate $p\in\mathcal{P}$ specifies either intrinsic properties of an object $o$ or relations between objects (e.g.,(cooked ?o1), (on ?o1 ?o2)). Each predicate is evaluated by a binary classifier over the continuous state, returning true or false. Grounding $\mathcal{P}$ across $\mathcal{S}$ produces a symbolic description of the environment $\mathcal{S}\times\mathcal{P}\rightarrow\mathcal{X}$. A logical action $a\in\mathcal{A}$ consists of a precondition $\mathcal{P}_{pre}$ $=\langle p_{1},p_{2},\dots\rangle$ and an effect $\mathcal{P}_{eff}=\langle p_{1}^{\prime},p_{2}^{\prime},\dots\rangle$. The precondition represents a set of predicates that must be satisfied for the action to be executed, while the effect describes the change of the resultant state upon action completion. Logical actions define the logical state transitions $\mathcal{X}^{(t)}\times a^{(t)}\rightarrow\mathcal{X}^{(t+1)}$. Thus, any planning problem $\langle\mathcal{S}^{(init)},\mathcal{S}^{(final)}\rangle$ can be formulated as a logical planning problem $\mathcal{Q}=\langle\mathcal{O},\mathcal{D},\mathcal{X}^{(init)},\mathcal{X}^{(final)}\rangle$, which is solved by a symbolic planner to produce a task plan ${a^{(0)},a^{(1)},\dots,a^{(T-1)}}=\texttt{PDDLSolver}(\mathcal{Q})$. Each logical action $a$ must then be integrated with corresponding motion planning skills to generate continuous robotic actions $\tilde{a}$ for execution.
@@ 4 Problem Statement
We address the robot planning problem given a single human demonstration $\tau_{demo}$ and its corresponding task description $T_{demo}$. The demonstration $\tau_{demo}$ is represented as a sequence of continuous environment states $\tau_{demo}=\{\mathcal{S}^{(0)},\mathcal{S}^{(1)},\dots\}$ and $T_{demo}$ is a brief natural language phrase. We assume the robot is the sole agent in the environment (Huang et al., 2024). Additionally, we assume that the demonstration covers the necessary domain knowledge to solve the target task. The proposed PDDLLM framework aims to generate a sequence of continuous robotic actions for a new planning problem. The problem can be formulated as:
$\tilde{a}^{(0)},\tilde{a}^{(1)},\dots,\tilde{a}^{(L-1)}=PDDLLM(\mathcal{S}_{new}^{(init)},\mathcal{S}_{new}^{(goal)},T_{demo},\tau_{demo})\,,$ (1)
where $\mathcal{S}_{new}^{(init)},\mathcal{S}_{new}^{(goal)}$ define the new planning problem and $L$ is the plan length.
@@ 5 Methodology
Figure 1 presents an overview of the PDDLLM framework. With $T_{demo}$ and $\tau_{demo}$ as inputs, PDDLLM constructs a relevant predicate library through predicate imagination and generates an action library via action invention. Ultimately, these predicate and action libraries are compiled into an executable PDDL planning domain, automatically interfaced with motion planners via LoCAS. In the following sections, we provide a detailed explanation of each step in the framework.
Figure 1: Overview of the proposed framework. (1) Human demonstrations, in the form of manipulation trajectories, and the corresponding task descriptions, serve as input. (2) PDDLLM initiates thousands of parallel simulations to imagine predicates by summarizing the outcomes with LLMs, returning the predicate library along with the relevance of each predicate to the current task. (3) Actions are invented by an LLM that summarizes logical state transition patterns from the demonstration, which is grounded into logical states using the imagined predicates. (4) The predicates and actions are compiled into a planning domain, which is automatically interfaced with motion planning algorithm by the Logic-Constrained Action Sampler (LoCAS) to solve new tasks.
@@ 5.1 PDDL Domain Generation
5.1.1 Predicate Imagination:
The process of predicate imagination consists of two stages. Stage (1) generate first-order predicates and Stage (2) futher expand the higher-order predicates.
Stage (1)
In this stage, PDDLLM generates first-order predicates, which directly describe the physical properties or relations of the objects (e.g., (is_on ?o1 ?o2), (smaller ?o1 ?o2)), by summarizing simulated object interactions using an LLM.
@@ 5.1.1 Predicate Imagination:
The feature space is defined as a set of continuous state variables, such as position, size, and color, that fully characterize the state of each object in the environment.
@@
Following Definition 1, the feature space is defined as a set of variables, such as Cartesian coordinates $(x,y,z)$ for object positions and RGB values $(r,g,b)$ for colors. The feature space is bounded by real-world constraints. Object states are uniformly sampled across this space to span a diverse range of object-object and object-environment interactions. Each sampled state undergoes a two-step verification process. First, it is evaluated for physical feasibility using a physical simulator, which serves as a physical knowledge base to capture complex dynamics beyond the reasoning capacity of LLMs. Only physically valid states are retained.
@@
Next, PDDLLM partitions the feature space into a finite collection of subspaces. The range of each feature is divided into intervals, with the length of each interval being a hyperparameter. The intersections of these intervals specify the subspaces. Each object state can be mapped into one of the subspaces. Each subspace is analyzed to determine whether it contains feasible object states, as verified through simulation in the previous verification step. If so, prompts are generated using a predefined template to describe the corresponding scenes. Prompt generation is automated as it only requires the replacement of some keywords (such as “position” with “color”) and the specification of interval boundaries. An LLM is then prompted to summarize subspaces into meaningful predicates and select those relevant to the task. The subspace boundaries serve as predicate implications, enabling the classification of whether a predicate holds true. Figure 2.a illustrates an example of predicate generation for positional relations between objects.
@@
Figure 2: a. This example illustrates the imagination of predicates for relative object positions. Let $u$ be a configurable variable determined by the feature subspace size. Object poses are sampled and simulated, with infeasible cases (e.g., Simulation 1) filtered out. Feasible subspaces are encoded into LLM prompts, and the LLM returns task-relevant predicates. b. This example shows how Stack actions are invented. Continuous states are grounded into logical states using the imagined predicates, where the state transition represents the logical action. By prompting the LLM with the transition from the initial state to the next state, we obtain the PDDL definition of action Stack. c. The integration of actions with the motion planner is handled automatically by LoCAS, which converts first-order predicates in the action effect $\mathcal{P}_{eff}$ into geometric constraints for motion planning. Prompt examples for a and b are provided in Section A.3.
Stage (2)
[…]
5.1.2 Action Invention:
@@
After constructing the predicate library, the human demonstration $\mathbf{\tau}_{demo}$ can be mapped into the logical space as $\mathbf{\tau}_{demo}^{logic}$ by grounding all relevant predicates at each time step. The logical state transitions within $\mathbf{\tau}_{demo}^{logic}$ signify the execution of actions. An advantage of learning actions in the logical space is that it simplifies pattern recognition by focusing on moments of logical state transitions, effectively transforming long trajectories into concise logical representations. For instance, the continuous manipulation trajectory $\mathbf{\tau}_{demo}$ in Figure 2.b contains over 1000 time steps while $\mathbf{\tau}_{demo}^{logic}$ is reduced to merely 2 steps. Logical state transitions are extracted from pre-and-post-change states and presented to the LLM as prompts. After experimenting with various prompt structures, we found this direct, structured method to be the most effective. A concrete example of action invention for Stack is elaborated in Figure 2.b.
5.2 Logic-Constrained Action Sampler (LoCAS)
Having obtained the logical action definitions, the next step is to establish an interface between each logical action and the low-level motion planner. This is achieved through the Logic-Constrained Action Sampler (LoCAS). Traditional motion planning algorithms search for robot pose sequences within the workspace until a feasible trajectory is found (Gammell et al., 2015). Later, Toussaint (2015) proposed encoding logical constraints as mathematical expressions for optimization. Combining these ideas with domain generation motivate us to develop LoCAS. It ensures that the first-order predicates in a logical action’s effect $\mathcal{P}_{\text{eff}}$ are satisfied upon action completion by enforcing them as constraints during the motion planning process, as is in Figure 2.c. These predicate implications, expressed as mathematical inequalities, guide the motion planner. Consequently, each logical action is transformed into a standard constrained motion planning problem. The adoption of LoCAS automatically grounds logical actions with motion planners and thus eliminates the need for predefined skills, which are human-designed or deep-learned task-specific motion planning policies commonly assumed available in previous studies (Kumar et al., 2023; Liang et al., 2024).
Planning Pipeline: The user provides a single human demonstration $\tau_{demo}$ and its corresponding task description $T_{demo}$, based on which PDDLLM automatically constructs a planning domain $\mathcal{D}_{\mathrm{PDDLLM}}$ and integrates it with low-level motion planners, enabling direct deployment in robot systems. When a novel planning task $(\mathcal{S}^{(init)}_{new},\mathcal{S}^{(final)}_{new})$ is presented, the initial and goal states are first mapped to their corresponding logical representations using the invented predicates. For the formulated logical planning problem $\mathcal{Q}_{new}=(\mathcal{O},\mathcal{D}_{\mathrm{PDDLLM}},\mathcal{X}^{(init)}_{new},\mathcal{X}^{(final)}_{new})$, PDDLSolver is used to solve for the task plan, which is refined into continuous action sequence $\{\tilde{a}^{(0)},\tilde{a}^{(1)},\dots,\tilde{a}^{(L-1)}\}$ by LoCAS.

### v3

@@ Abstract
Pre-trained large language models (LLMs) show promise for robotic task planning but often struggle to guarantee correctness in long-horizon problems. Task and motion planning (TAMP) addresses this by grounding symbolic plans in low-level execution, yet it relies heavily on manually engineered planning domains. To improve long-horizon planning reliability and reduce human intervention, we present Planning Domain Derivation with LLMs (PDDLLM), a framework that automatically induces symbolic predicates and actions directly from demonstration trajectories by combining LLM reasoning with physical simulation roll-outs. Unlike prior domain-inference methods that rely on partially predefined or language descriptions of planning domains, PDDLLM constructs domains without manual domain initialization and automatically integrates them with motion planners to produce executable plans, enhancing long-horizon planning automation. Across 1,200 tasks in nine environments, PDDLLM outperforms six LLM-based planning baselines, achieving at least 20% higher success rates, reduced token costs, and successful deployment on multiple physical robot platforms.
@@ 1 Introduction
Robotic planning remains challenging in complex scenarios that require abstract, long-horizon reasoning. Large language models (LLMs) demonstrate strong generalization in this domain but often struggle with temporal dependencies in extended tasks Huang et al. (2022, 2023). Task and Motion Planning (TAMP) frameworks provide robustness in long-horizon reasoning by integrating high-level symbolic reasoning with low-level motion planning. However, they face two major limitations: (i) planning domains, expressed in symbolic planning languages such as PDDL (McDermott et al., 1998), are difficult to ground in complex geometric information, and (ii) these domains are labor-intensive to manually construct Garrett et al. (2020, 2021); Khodeir et al. (2023); Silver et al. (2020). Recent vision-language-action models have advanced the first challenge by improving geometric grounding through semantic action instructions, but they pay limited attention to long-horizon reasoning Black et al. (2024); Kim et al. (2024); Zitkovich (2023); Li et al. (2024); Team (2024). Complementarily, our work focuses on solving the second challenge by reducing the human effort required for domain construction, thereby enabling automated and scalable long-horizon reasoning.
Existing domain-generation methods typically rely on predesigned elements (predicates or actions) to complete a domain Silver et al. (2023); Kumar et al. (2023), where manual predesign remains critical and often depends on curated training data Huang et al. (2025). LLM-based approaches also demand extensive natural-language descriptions of PDDL domains and careful prompt engineering Guan et al. (2023a); Oswald et al. (2025). Meanwhile, recent advances in world-model learning show that LLMs can serve as compact explainers and reasoners for robot behavior and object relations (Guan et al., 2023b; Zhao et al., 2024; Liang et al., 2024; Tang et al., 2024). Motivated by this, we propose to combine LLM reasoning with physical simulation to generate planning domains, eliminating reliance on manual predesign, curated training data, and detailed textual descriptions.
We introduce PDDLLM, the first method to construct planning domains without manual predesign, training data, or extensive human descriptions. Moreover, PDDLLM automates integration with low-level motion planners, further reducing the need for manual intervention. Our method targets at long-horizon planning problems, aligning with the scope of TAMP studies. Managing action sequences and task structures is essential for these tasks such as household activities and puzzle solving. The main contributions of this paper are:
  • An algorithm that combines LLMs and physical simulation to automatically generate a human-interpretable planning domain from a single demonstration.
  • A Logical Constraint Adapter (LoCA) that systematically interfaces the generated domain with low-level motion planners.
  • An extensive evaluation on over 1,200 tasks in nine environments, demonstrating superior long-horizon planning performance and more efficient token usage than state-of-the-art baselines, with successful deployment on three physical robot platforms.
@@ 2 Related Work
A recent line of research aims to infer the planning domain directly from human demonstrations, environment interactions, or natural language. However, these approaches often depend on partially or fully predefined symbolic predicates and actions as well as extensive training dataset (Silver et al., 2023; Kumar et al., 2023; Huang et al., 2025; Wong et al., 2023; Liu et al., 2025). Some recent studies have explored leveraging LMs for domain generation, primarily by extending manually defined domains with additional predicates and actions (Liang et al., 2024; Athalye et al., 2024; Byrnes et al., 2024). The approaches proposed by Guan et al. (2023b); Han et al. (2024) generates planning domains requiring prompts containing human-crafted planning examples or detailed descriptions of predicates and actions. Moreover, many of these studies assume that symbolic actions must be manually linked to existing motion-level skills (Kumar et al., 2023; Huang et al., 2025; Liang et al., 2024; Athalye et al., 2024), requiring labor-intensive alignment between planning domains and low-level motion planners. LoCA addresses this limitation by automatically grounding symbolic actions into motion constraints, thereby reducing manual effort and enhancing scalability and autonomy.
@@ 3 Preliminaries
PDDL is a standard formal language used to specify planning problems. The object set $\mathcal{O}$ represents the environment’s objects, whose continuous state $\mathcal{S}$, such as pose, color, and size, can be queried via a perception function $\mathcal{I}:\mathcal{O}\times\mathcal{I}\rightarrow\mathcal{S}$. The PDDL domain $\mathcal{D}=(\mathcal{P},\mathcal{A})$ describes the general rules of the environment, consisting of a set of logical predicates $\mathcal{P}$ and a set of logical actions $\mathcal{A}$. A logical predicate $p\in\mathcal{P}$ specifies either intrinsic properties of an object $o$ or relations between objects (e.g.,(cooked ?o1), (on ?o1 ?o2)). Each predicate is evaluated by a binary classifier over the continuous state, returning true or false. A symbolic descripton of the environment $\mathcal{X}$ can be obainted by grounding $\mathcal{P}$ across $\mathcal{S}$ (i.e. $\mathcal{S}\times\mathcal{P}\rightarrow\mathcal{X}$). A logical action $a\in\mathcal{A}$ consists of a precondition $\mathcal{P}_{pre}$ $\langle p_{1},p_{2},\dots\rangle$ and an action’s effect $\mathcal{P}_{eff}=\langle p_{1}^{\prime},p_{2}^{\prime},\dots\rangle$. The precondition represents a set of predicates that must be satisfied for the action to be executed, while the action’s effect describes the change of the resultant state upon action completion. Logical actions define the logical state transitions $\mathcal{X}^{(t)}\times a^{(t)}\rightarrow\mathcal{X}^{(t+1)}$. Thus, any planning problem $\langle\mathcal{S}^{(init)},\mathcal{S}^{(final)}\rangle$ can be formulated as a logical planning problem $\mathcal{Q}=\langle\mathcal{O},\mathcal{D},\mathcal{X}^{(init)},\mathcal{X}^{(final)}\rangle$, which is solved by a symbolic planner to produce a task plan ${a^{(0)},a^{(1)},\dots,a^{(T-1)}}=\texttt{PDDLSolver}(\mathcal{Q})$. Each logical action $a$ must then be integrated with corresponding motion planning skills to generate continuous robotic actions $\tilde{a}$ for execution.
@@ 4 Problem Statement
Our system solves a new long-horizon planning problem given an expert demonstration trajectory $\tau_{demo}$ and its associated task description $T_{demo}$. The trajectory $\tau_{demo}$ is a sequence of continuous environment states $\mathcal{S}^{(0)},\mathcal{S}^{(1)},\dots$}, while $T_{demo}$ is a short natural language phrase describing the task. Following prior work (Silver et al., 2023; Kumar et al., 2023; Huang et al., 2025), the robot is the sole acting agent during data collection.
Given a new planning problem, PDDLLM derives a PDDL domain from $(T_{demo},\tau_{demo})$ which generates a sequence of symbolic actions as the task plan. LoCA, a component within PDDLLM, interfaces the task plan with an off-the-shelf motion planner, which produces executable robot trajectories. Formally, the problem can be expressed as:
$\tilde{a}^{(0)},\tilde{a}^{(1)},\dots,\tilde{a}^{(L-1)}$ $=MotionPlanner(PDDLLM(\mathcal{S}^{(init)}_{new},\mathcal{S}^{(goal)}_{new},T_{demo},\tau_{demo}))$ (1)
where $\mathcal{S}_{new}^{(init)}$ and $\mathcal{S}_{new}^{(goal)}$ define the initial and goal states of the new problem, $L$ denotes the task plan length, and $\tilde{a}^{(0:L)}$ indicates an executable robot trajectory.
@@ 5 Methodology
Figure˜1 presents an overview of the PDDLLM framework. With $T_{demo}$ and $\tau_{demo}$ as inputs, PDDLLM constructs a relevant predicate library through predicate imagination and generates an action library via action invention. Ultimately, these predicate and action libraries are compiled into an executable PDDL planning domain, automatically interfaced with motion planners via LoCA. In the following sections, we provide a detailed explanation of each step in the framework.
Figure 1: Overview of the proposed framework. (1) Human demonstrations, in the form of manipulation trajectories, and the corresponding task descriptions, serve as input. Implementation details is shown in Section˜A.12. (2) PDDLLM initiates thousands of parallel simulations to imagine predicates by summarizing the simulation roll-outs with LLMs, returning the predicate library along with the relevance of each predicate to the current task. (3) Actions are invented by an LLM that summarizes logical state transition patterns from the demonstration, which is grounded into logical states using the imagined predicates. (4) The predicates and actions are compiled into a planning domain, which is automatically interfaced with motion planning algorithm by the Logical Constrain Adapter (LoCA) to solve new tasks.
@@ 5.1 PDDL Domain Generation
5.1.1 Predicate Imagination
Predicate imagination refers to the process of summarizing simulated object relation roll-outs into meaningful predicates. The process consists of two stages: Stage 1 generates first-order predicates and Stage 2 further derives the higher-order predicates from the first-order predicates.
Stage 1. In this stage, PDDLLM generates first-order predicates, which directly describe the physical properties or relations of the objects (e.g., (is_on ?o1 ?o2), (smaller ?o1 ?o2)), by summarizing simulated object interactions using an LLM.
@@ 5.1.1 Predicate Imagination
The feature space is defined as a set of continuous state variables, such as position, orientation, size, and color, that fully characterize the state of each object in the environment.
@@
Following Definition˜1, the feature space is defined as a set of variables, such as Cartesian coordinates $(x,y,z)$ for object positions and RGB values $(r,g,b)$ for colors. The feature space is bounded by real-world constraints. Object states are uniformly sampled across this space to span a diverse range of object-object and object-environment interactions. Each sampled state undergoes a two-step verification process. First, it is evaluated for physical feasibility using a physical simulator, which serves as a physical knowledge base to capture complex dynamics beyond the reasoning capacity of LLMs. Only physically valid states are retained.
@@
Next, PDDLLM partitions the feature space into a finite collection of subspaces. The range of each feature is divided into intervals, with the length of each interval being a hyperparameter. The intersections of these intervals specify the subspaces. Each object state can be mapped into one of the subspaces. Each subspace is analyzed to determine whether it contains feasible object states, as verified through simulation in the previous verification step. If so, simulation roll-out summaries are generated as prompts following a predefined scene-description template. Prompt generation is automated as it only requires the replacement of some keywords (such as “position” with “color”) and the specification of interval boundaries. An LLM is then prompted to summarize subspaces into meaningful predicates and select those relevant to the task. The subspace boundaries serve as predicate physical constraints, enabling the classification of whether a predicate holds true. Figure˜2.a illustrates an example of predicate generation for positional relations between objects.
@@
Figure 2: a. This example illustrates the imagination of predicates for relative object positions. Let $u$ be a configurable variable determined by the feature subspace size. Object poses are sampled and simulated, with infeasible cases filtered out. Feasible subspaces are provided to the LLM to generate first-order predicates with their corresponding physical constraints. Higher-order predicates can be further derived using logical operators (e.g., "not", "for all") from first-order predicates. Diverse predicate examples are provided in Section˜A.9 b. This example shows how the Stack action is invented. Continuous states are grounded into logical states using the imagined predicates, where the state transition represents the logical action. By prompting the LLM with the pair of the current state and the next state after the transition, we obtain the PDDL definition of the action Stack. c. The integration of actions with the motion planner is handled automatically by LoCA, which retrieves the physical constraints associated with each first-order predicate in the action effect set $\mathcal{P}_{eff}$ and applies these constraints for motion planning. Prompt examples for a and b are provided in Section˜A.3.
Stage 2. First-order predicates are combined with logical operators and quantifiers to construct higher-order predicates. They are systematically combined in all possible ways to generate more complex logical expressions. Following prior work (Curtis et al., 2022; Silver et al., 2023), we primarily use the negation operator, along with the quantifiers “for all” ($\forall$) and “there exists” ($\exists$). For example, (is_on ?o1, ?o2) can be negated to produce (not_is_on ?o1, ?o2). When combined with the universal quantifier, this further yields ($\mathrm{}{\forall}$_o1_not_is_on ?o1, ?o2), meaning that for any o1 in the environment, (is_on ?o1, ?o2) is false. This indicates that object o2 is on top.
5.1.2 Action Invention
@@
After constructing the predicate library, the human demonstration $\mathbf{\tau}_{demo}$ can be mapped into the logical space as $\mathbf{\tau}_{demo}^{logic}$ by grounding all relevant predicates at each time step. The logical state transitions within $\mathbf{\tau}_{demo}^{logic}$ signify the execution of actions. An advantage of learning actions in the logical space is that it simplifies pattern recognition by focusing on moments of logical state transitions, effectively transforming long trajectories into concise logical representations. For instance, the continuous manipulation trajectory $\mathbf{\tau}_{demo}$ in Figure˜2.b contains over 1000 time steps while $\mathbf{\tau}_{demo}^{logic}$ is reduced to merely 2 steps. Logical state transitions are extracted from pre-and-post-change states and presented to the LLM as prompts. After experimenting with various prompt structures, we found this direct, structured method to be the most effective. A concrete example of action invention for Stack is elaborated in Figure˜2.b.
5.1.3 Generation Refinement
To improve the quality and accuracy of the generated planning domains, we adopt a parallel prompting strategy. Specifically, multiple candidate PDDL models are generated from the same demonstration, and candidates that fail to produce feasible task plans are replaced by alternatives. Empirically, we find that when the number of parallel prompts exceeds five, the correctness of the final domain converges to a stable level. Examples of the parallel prompting procedure are shown in Section˜A.7
5.2 Automatic Interfacing with Motion Planner
Combining the PDDL planning domain with symbolic solvers yields a task plan $a^{(0)},a^{(1)},\dots,a^{(T-1)}$}. The next step is to establish an interface between each logical action and the low-level motion planner.
In PDDLLM, this interfacing is fully automated through the Logical Constraint Adapter (LoCA). Prior approaches relied on manually encoding logical actions as mathematical constraints for motion planning (Toussaint, 2015). LoCA automates and generalizes this process by directly retrieving the physical constraints associated with the first-order predicates in an action’s effect set $\mathcal{P}_{\text{eff}}$. These constraints, expressed as mathematical inequalities, are sequentially applied to the motion planner, ensuring that the required predicates are satisfied once the action is executed. In this way, LoCA automatically transforms each logical action into a standard constrained motion planning problem. As a result, the generated trajectory is guaranteed to align with the semantic meaning of the action, as illustrated in Figure˜2.c. This eliminates the need for manually engineered interfaces or predefined motion-planning skills, thereby streamlining the integration between symbolic planning and low-level control.
Additionally, our framework is also compatible with recent advances such as vision–language–action (VLA) models by directly using logical actions as prompts. Experiments in Section˜8 demonstrate PDDLLM’s ability to automatically integrate with both traditional and learning-based motion planning.